OpenStudy (anonymous):

solve by the substitution method 15x-10y=5 5y= -3+10x

OpenStudy (anonymous):

15x-2(5y)=5 => 15x-2(-3+10x)=5 =>15x+6-20x=5 =>-5x=-1=>x=1/5 5y=-3+10x=-3+2=-1=>y=-1/5